WARNING: This web is not used anymore. Please use PDBService.MinuteS11January07 instead!
 

Minutes 11 January 07

  • Phone:
    • TRIUMF: Denise
    • GridKA: Andreas, Doris
    • FNAL: Eric
    • CNAF: Barbara
    • ATLAS: Shasha, David Front
    • SARA: Alexander
    • BNL: Iris
    • NDGF: Olli

  • CERN:
    • 3D: Dirk, Zbigniew, Marcin, Luis, Eva
    • RAL: Gordon
    • ATLAS: Gancho, Florbela

  • General announcements:
    • 3D status report by Dirk at the GDB meeting
    • next workshop 22 - 26 January

  • Sites update:
    • Triumf (Denice):
      • RAC working well on one node, network/memory problems on one of the nodes
      • working on the problem, Denice will mail the general list when the problem is solved
    • CNAF (Barbara):
      • installation next week
      • in production before half/end February
      • tests on the LFC replica to be planned
      • Barbara asks if it is possible to include the master LFC (LHCB production database) in the Streams monitoring - to be checked
    • BNL (Iris):
      • memory to be added (already delivered)
      • patch set will be applied after increasing the memory
      • waiting for metalink account
        • Dirk will collect all the information and will send an email with the missing steps, and will create the metalink accounts for the sites
    • GridKA (Andreas):
      • upgrade to 10.2.0.3 forseen 21st January
        • cursor problem discovered in this version - additional patch to be installed
        • streams performance impacted (up to 10%)
    • SARA:
      • database ready ( 2 nodes cluster)
      • already included on the Streams monitoring
      • Florbela pointed out that the streams pool size is too small and there is no info available using oem - Eva will check with Alexander
    • NDGF (Olli):
      • single node database will be ready soon
      • ATLAS will require 3 nodes database - check if it is possible to fulfill the requirements
    • RAL (Gordon):
      • monitoring mail list to be provided - Zbigniew will send a remainder

  • Agenda fro the workshop:
    • Dirk has sent the agenda to the full list
    • Gordon will collect all the information related to Tier1 sites problems
    • Barbara will report on the LFC Streams replication

  • ATLAS Streams tests:
    • first successful run: 4 sites during 12 hours
    • evolution of the storage to be discussed during the workshop ( 1TB of data, and 2-3 TB of additional structures: indexes, ... to allow TAGS querying)
    • conditions data challenge postponed - end february??
    • SARA seems to keep the rate (first tests)
    • Triumf and BNL: network performance problems - to be checked with the networking people
Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2007-02-01 - EvaDafonte
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback